# for Large Language Models

### v2

@@ 4.1.4. Implementation Details
For our model implementation, we primarily utilize the PyTorch and Transformers libraries. We employ Vicuna-7B-v1.1 and Vicuna-7B-v1.5 as the base models for our approach. The batch size is set to 2 on each GPU, and the learning rate is $2e^{-3}$. We apply a warmup ratio of $3e^{-2}$ and set the maximum input length of the Large Language Model (LLM) to 2048. The training process is carried out for 3 epochs. In the stage of task-specific instruction tuning, we explore the performance of the model under different data mixtures by adopting various combinations of instruction data. The hyperparameter settings remain constant, except for the number of training epochs, which is set to 2 in this stage. The alignment projector parameters fine-tuned in the self-supervised instruction tuning stage are used as the initial parameters for the projector in the second tuning stage. For the evaluation of most baselines, we utilize their publicly available code. We employ a grid-search strategy based on their default hyperparameter settings to ensure a comprehensive evaluation. For further implementation details, please refer to our released source code.
@@ 4.2. Overall Performance Comparison (RQ1)
We conduct experiments on the node classification task, evaluating both supervised and zero-shot scenarios. The overall performance is presented in Table 1. Supervised Task Settings: We train the models on a specific dataset and evaluated their performance on the corresponding test set (e.g., training on Arxiv-Arxiv and testing on the Arxiv test set). Zero-Shot Task Settings: We train the models on a specific dataset and test them on other datasets without any additional training (e.g., training on Arxiv-PubMed and testing on the PubMed dataset). To account for variations in the number of classes across different datasets, we employed a classifier trained with transfer data, typically a linear layer, when testing GNN-based models. In Table 1, ”-7B-” represents the parameter scale, while ”-v1.1-” and ”-v1.5-” indicate different versions of the base Vicuna model. ”-stage2” indicates that only the second stage tuning is adopted. ”-std” and ”-cot” denote the use of the standard and generated COT instruction datasets, respectively.
@@
Additionally, LLM-based solutions, like Baichuan-7B and Vicuna-7B maintain stable performance across different datasets. However, they are limited to making predictions solely based on text information. In contrast, our GraphGPT effectively preserves graph structured information, offers a more comprehensive solution for graph learning tasks. These improvements can be attributed to two key factors: i) Through our dual-stage graph instruction tuning, our method aligns the graph tokens, which contain rich structural information encoded by the graph encoder, with the natural language tokens. This alignment allows the LLM to retain and understand the inherent structural characteristics of the graph data. ii) Our framework facilitates mutual enhancement between the graph encoder and LLM. The introduction of graph tokens fills the gap in the LLM’s structural understanding, enabling it to incorporate and reason about the graph’s structural information.
@@ 4.3. Generalization Ability Investigation (RQ2)
Generalization for Multitasking Graph Learner. Recent studies on instruction tuning suggest that mixing different instruction tuning data can further enhance the performance of Language and Logic Models (LLMs). In this study, we ensure a consistent number of instruction entries and mix different types of instruction data, including standard instruction (”-std”), COT instruction (”-cot”), a blend of standard (50%) and COT (50%) instruction (”-mix”), and link prediction instruction (”Link”). The results are presented in Tables 2 and Table 7 in Appendix. We can observe that effective data mixture solutions can significantly improve the performance of our GraphGPT under various settings. The addition of task-specific instruction for link prediction task notably enhances the performance of our model in node classification. Interestingly, after incorporating node classification, the performance of link prediction also exceeds that of the selected best-performing existing models. After mixing the instructions of different tasks, our model demonstrates the ability to effectively handle various graph learning tasks and transfer its knowledge to other unseen datasets.
@@ 4.6. Model Case Study (RQ5)
We conduct a detailed analysis of our model’s performance in downstream graph learning tasks and compare it to traditional LLMs using different types of instructions. We prompt both ChatGPT and our GraphGPT with Arxiv data, using only node content (title and abstract), node content with text-based graph structure, and our designed graph instruction. The results, presented in Table 5 and 10 in the appendix, clearly demonstrate that ChatGPT, despite its massive parameter count (over 200B), struggles to make accurate predictions based solely on node text information or node content with the text-based graph structure. This difficulty is particularly pronounced when dealing with papers that have a high degree of cross-disciplinary characteristics, as illustrated in the example of machine learning and hardware architecture. In contrast, our GraphGPT consistently provides accurate predictions and offers reasonable explanations. This is because our GraphGPT accepts a subgraph structure with 103 nodes, allowing it to extract rich structural information from the neighboring nodes’ citation relationships, enabling accurate predictions.
@@ 5. Related Work
Self-supervised Learning and Pre-training on Graphs. To enhance the robustness of graph models, self-supervised learning (SSL) has been introduced as a powerful technique (Liu et al., 2022; Jing et al., 2021; Hu et al., 2020b). It allows GNNs to learn meaningful graph representations without heavily relying on labeled data. The core idea behind self-supervised learning in graph models is to design pretext tasks that leverage the graph’s intrinsic properties to generate additional supervision signals (Xia et al., 2022). SSL-enhanced graph learning methods can be broadly classified into two main paradigms: contrastive SSL and generative SSL. In particular, i) Contrastive SSL focuses on learning representations by contrasting positive and negative samples. Notable methods in this domain include GraphCL (You et al., 2020) and GCA (Zhu et al., 2021). Recent advancements in contrastive SSL include automated contrastive augmentation (i.e., JOAO (You et al., 2021), AutoGCL (Yin et al., 2022)), Hyperbolic-Euclidean due space contrasting (e.g., DSGC (Yang et al., 2022a)), or community-aware contrastive learning (e.g., gCooL (Li et al., 2022)). ii) Generative SSL, on the other hand, focuses on generating realistic samples that resemble the original graph structures. Recent advancements in this line include GraphMAE (Hou et al., 2022, 2023) for feature masking, and S2GAE (Tan et al., 2023), AutoCF (Xia et al., 2023) for reconstructing masked edges as SSL tasks.
@@ 6. Conclusion
Appendix A Appendix
In the appendix, we provide a summary of the statistical information of our datasets in Section A.1. Furthermore, in Section A.2, we offer detailed descriptions of the baseline methods used for comparison in our experiments. To facilitate the understanding of the instruction templates used for different tasks and datasets, we present them in Section A.4. These templates serve as guidelines for formulating instructions that are specific to the task and dataset at hand. In Section A.3, we present the implementation details of the text-graph grounding process. Finally, in Section A.5, we provide supplementary experimental results.
A.1. Statistical information of datasets
The detailed statistical information of datasets is shown in Table 6.
Table 6. Statistical information of datasets.
Dataset # Nodes # Edges # Classes
ogb-arxiv 169,343 1,166,243 40
PubMed 19,717 44,338 3
Cora 25,120 182,280 70
Table 7. Performance comparison of various instruction mixtures for link prediction on the PubMed dataset.
Dataset PubMed
Model AUC AP
MLP 0.5583 0.5833
GAT 0.5606 0.6373
GraphSAGE 0.5041 0.5813
RevGNN 0.4538 0.5083
Node2Vec 0.6535 0.6885
w/o Link 0.5010 0.5005
only Link 0.6704 0.6087
Arxiv-std + PubMed-std + Link 0.8246 0.8026
Arxiv-mix + PubMed-mix + Link 0.6451 0.5886
Table 8. Instruction template for the graph matching task.
Graph Matching
Input: Given a sequence of graph tokens <graph> that constitute a subgraph of a citation graph, where the first token represents the central node of the subgraph, and the remaining nodes represent the first and second order neighbors of the central node. Each graph token contains the title and abstract information of the paper at this node. Here is a list of paper titles: {1. Paper 1. 2. Paper 2. ... n. Paper n}, please reorder the list of papers according to the order of graph tokens (i.e., complete the matching of graph tokens and papers).
Output: Based on the given graph tokens and the list of paper titles, we obtain the matching of graph tokens and papers as follows: {Graph token 1 corresponds to Paper Q, Graph token 2 corresponds to Paper P, ... Graph token n corresponds to Paper R}.
Table 9. Instruction template for node classification and link prediction on different datasets.
Node Classification (Arxiv in the COT manner)
Input: Given a citation graph: <graph> where the 0th node is the target paper, with the following information: Abstract: {abstract of the central node}. Title: {title of the central node}. Question: Which arXiv CS sub-category does this paper belong to? Give 5 likely arXiv CS sub-categories as a comma-separated list ordered from most to least likely, in the form ”cs.XX”. Please think about the categorization in a step by step manner and avoid making false associations. Then provide your reasoning.
Output: Based on the information, {reasoning process and answers}.
Arxiv in the standard manner)
Input: Given a citation graph: <graph> where the 0th node is the target paper, and other nodes are its one-hop or multi-hop neighbors, with the following information: Abstract: {abstract of the central node}. Title: {title of the central node}. Question: Which arXiv CS sub-category does this paper belong to? Give the most likely arXiv CS sub-categories of this paper directly, in the form ”cs.XX” with full name of the category.
Output: {the ground-truth answer}.
Node Classification (PubMed in the COT manner)
Input: Given a citation graph: <graph> where the 0th node is the target paper, with the following information: Abstract: {abstract of the central node}. Title: {title of the central node}. Question: Does the paper involve any cases of Type 1 diabetes, Type 2 diabetes, or Experimentally induced diabetes? Please give one or more answers of either Type 1 diabetes, Type 2 diabetes, or Experimentally induced diabetes; if multiple options apply, provide a comma-separated list ordered from most to least related. Please think about the categorization in a step by step manner and avoid making false associations. Then provide your reasoning for each choice.
Output: Based on the information, {reasoning process and answers}.
Node Classification (PubMed in the standard manner)
Input: Given a citation graph: <graph> where the 0th node is the target paper, and other nodes are its one-hop or multi-hop neighbors, with the following information: Abstract: {abstract of the central node}. Title: {title of the central node}. Question: Which case of Type 1 diabetes, Type 2 diabetes, or Experimentally induced diabetes does this paper involve? Please give one answer of either Type 1 diabetes, Type 2 diabetes, or Experimentally induced diabetes directly.
Output: {the ground-truth answer}.
Node Classification (Cora in the COT manner)
Input: Given a citation graph: <graph> where the 0th node is the target paper, with the following information: Abstract: {abstract of the central node}. Title: {title of the central node}. Question: Which of the following subcategories of computer science does this paper belong to: {1. Categories 1. 2. Categories 2. ... n. Categories n}? Give 5 likely categories as a comma-separated list ordered from most to least likely. Please think about the categorization in a step by step manner and avoid making false associations. Then provide your reasoning for each choice.
Output: Based on the information, {reasoning process and answers}.
Node Classification (Cora in the standard manner)
Input: Given a citation graph: <graph> where the 0th node is the target paper, with the following information: Abstract: {abstract of the central node}. Title: {title of the central node}. Question: Which of the following subcategories of computer science does this paper belong to: {1. Categories 1. 2. Categories 2. ... n. Categories n}? Directly give the full name of the most likely category of this paper.
Output: {the ground-truth answer}.
Link Prediction (PubMed)
Input: Given a sequence of graph tokens: <graph> that constitute a subgraph of a citation graph, where the first token represents the central node of the subgraph, and the remaining nodes represent the first and second order neighbors of the central node. The information of the central node is as follow: Abstract: {abstract of the central node}. Title: {title of the central node}. The other sequence of graph tokens: <graph>, where the first token (the central node) with the following information: Abstract: {abstract of the central node}. Title: {title of the central node}. If the connections between nodes represent the citation relationships between papers, are these two central nodes connected? Give me a direct answer of ”yes” or ”no”.
Output: {the ground-truth answer}.
A.2. Detailed Descriptions of Baselines
For a comprehensive performance comparison, our baseline set consists of the following methods which are presented below:
1) Conventional Deep Learning Methods
  • MLP: This method employs a multilayer perceptron to independently encode different node representations.
2) Graph Neural Encoders
  • GraphSAGE (Hamilton et al., 2017): It is a framework designed for inductive representation learning on large graphs, enabling the efficient generation of node embeddings for previously unseen data.
  • GCN (Kipf and Welling, 2017): This method extends the solution of convolutional neural networks to model graph-structured features.
  • GAT (Velickovic et al., 2018): This method introduces graph attention networks that leverage masked self-attentional layers to address the limitations of GCN. It achieves this by differentiating the message passing among different nodes with weighted information aggregation.
  • RevGNN (Li et al., 2021): This work investigates several techniques, including reversible connections, group convolutions, weight tying, and equilibrium models, to enhance the memory and parameter efficiency of graph neural networks (GNNs).
3) Self-Supervised Graph Learning Approaches
  • DGI (Velickovic et al., 2019): It maximizes the mutual information between patch representations and corresponding high-level summaries of graphs. By doing so, DGI enables the model to capture meaningful information from local patches of the graph and leverage it to learn informative node representations in an unsupervised manner.
4) Graph Knowledge Distillation Frameworks
  • GKD (Yang et al., 2022b): This method distills knowledge from a teacher GNN trained on a complete graph to a student GNN operating on a smaller or sparser graph. It achieves this by utilizing the Neural Heat Kernel (NHK) method, which encapsulates the geometric property of the underlying manifold.
  • GLNN (Zhang et al., 2022b): This work combines the advantages of graph neural networks and MLPs using knowledge distillation, with the aim of removing the dependency on the inference graph.
5) Graph Transformer Networks
  • NodeFormer (Wu et al., 2023b): This work leverages a kernelized Gumbel-Softmax operator to reduce the computational complexity of the graph transformer model, while also enabling the learning of latent graph structures from large graphs.
  • DIFFormer (Wu et al., 2023a): This work introduces an energy-constrained diffusion-based graph transformer. This transformer is designed to encode a batch of instances into evolutionary states by incorporating information from other instances.
6) Large Language Models
  • Baichuan-7B (Yang et al., 2023): It is an open-source, large-scale pre-trained model with 7 billion parameters. It is specifically designed to be bilingual, supporting both Chinese and English languages. This model has been trained extensively on a diverse range of data.
  • vicuna-7B-v1.1 (Chiang et al., 2023): It is an open-source chatbot that has undergone fine-tuning using user-shared conversations collected from ShareGPT. The base model used for this fine-tuning is Llama-1, a reference to a specific model architecture.
  • vicuna-7B-v1.5 (Chiang et al., 2023): This model is an enhanced iteration of vicuna-7b-v1.1, building upon Llama-2 as its base model.
Table 10. Comparison of our GraphGPT’s predictions with ChatGPT on a arxiv paper (Continued).
Title: Greening Cloud-Enabled Big Data Storage Forensics: Syncany as a Case Study. (with 41 first-order and second-order neighbors)
Ground-Truth Category: cs.CR, Cryptography and Security
ChatGPT (Node Content Only): Based on the title and abstract of the paper, we can make an educated guess about the likely arXiv CS sub-categories that this paper belongs to. Here is a step-by-step breakdown of the reasoning: 1. The paper mentions ”cloud-enabled big data storage solutions” as the focus of investigation. This suggests that the paper is related to cloud computing and big data. 2. The paper also mentions ”cyber forensics” and ”investigation process” in the context of cloud-enabled big data platforms. This indicates that the paper is specifically about the forensic analysis of data stored in cloud-enabled big data platforms. Based on these observations, here are five likely arXiv CS sub-categories for this paper: 1. cs.DC (Distributed, Parallel, and Cluster Computing) - This sub-category covers topics related to distributed computing systems, which includes cloud computing. 2. cs.DB (Databases) - This sub-category covers topics related to database systems, which are often used in big data storage solutions. 3. cs.CR (Cryptography and Security) - This sub-category covers topics related to cyber security, which includes forensic analysis of digital evidence. 4. cs.SE (Software Engineering) - This sub-category covers topics related to software development and engineering, which may include the development of cloud-enabled big data storage solutions. 5. cs.CY (Computers and Society) - This sub-category covers topics related to the impact of computing on society, which may include the ethical and legal implications of forensic analysis in cloud-enabled big data storage.
ChatGPT (Node Content and Text-based Graph Structure): Based on the title of the paper ”Greening Cloud Enabled Big Data Storage Forensics: Syncany as a Case Study” we can make an initial assumption that the paper belongs to the ”Information Retrieval” (cs.IR) sub-category. However, to determine the most likely arXiv CS sub-categories, we need to analyze the content and context of the paper. 1. cs.DB (Databases): The paper focuses on cloud-enabled big data storage, which is closely related to databases. It discusses the identification, collection, analysis, preservation, and archiving of digital evidences in the context of Syncany private cloud storage service. This suggests that the paper may belong to the cs.DB sub-category. 2. cs.CR (Cryptography and Security): The paper addresses the forensic investigation of cloud-enabled big data storage platforms, which involves analyzing and recovering residual artifacts of criminal activities. This indicates a strong connection to the field of cybersecurity and digital forensics, which falls under the cs.CR sub-category. 3. cs.SE (Software Engineering): The paper may also belong to the cs.SE sub-category as it discusses the investigation process and the challenges involved in analyzing cloud-enabled big data storage platforms. It explores techniques and methodologies for locating and documenting residual artifacts of forensic value, which are relevant to the field of software engineering. 4. cs.CY (Computers and Society): The paper touches upon the societal impact of cloud-enabled big data storage and the need for efficient forensic investigation techniques. It highlights the potential reduction in time and resources spent on real-world investigations, which aligns with the concerns and issues addressed in the cs.CY sub-category. 5. cs.DC (Distributed, Parallel, and Cluster Computing): The paper mentions the decentralization of digital evidence acquisition in the context of BitTorrent Sync. This suggests a connection to distributed computing and parallel processing, which are topics covered in the cs.DC sub-category. Based on the above analysis, the likely arXiv CS sub-categories for the paper, ordered from most to least likely, are: cs.DB, cs.CR, cs.SE, cs.CY, cs.DC.
GraphGPT: Based on the title and abstract of the paper, we can infer that the paper is related to cyber forensics and cloud-enabled big data storage. [Uncaptioned image] We can then proceed to categorize it based on the sub-categories of arXiv CS. 1. cs.CR: This category is for cryptography research, which involves the study of cryptographic algorithms, protocols, and systems. While the paper does not explicitly mention cryptography, it does involve the analysis of data remnants on a cloud-enabled big data platform, which could potentially involve cryptographic techniques. 2. cs.DS: This category is for database systems, which involves the design, implementation, and management of databases. The paper mentions the use of a popular storage engine for big data platforms, which suggests that the paper could be related to database systems. 3. cs.IR: This category is for information retrieval, which involves the study of how to effectively retrieve and organize information from large collections of data. The paper’s focus on identifying and documenting residual artifacts of forensic value on a cloud-enabled big data platform suggests that it could be related to information retrieval. 4. cs.IT: This category is for information technology, which involves the study of the design, development, and use of information technology systems. The paper’s focus on cloud-enabled big data storage and forensics suggests that it could be related to information technology. 5. cs.SY: This category is for systems and networking, which involves the study of computer systems and networks. The paper’s focus on the analysis of data remnants on a cloud-enabled big data platform suggests that it could be related to systems and networking.
A.3. Details of Text-Graph Grounding
For the implementation details of Text-Graph Grounding 3.1, we follow previous work (Wen and Fang, 2023), aligning text features and graph features across three dimensions. Mathematically, these dimensions can be represented by three sets of transformation equations:
$g_{1}^{(1)}(\hat{\mathbf{H}})=\{\hat{\mathbf{H}}_{i},1\leq i\leq N\}$ $,g_{1}^{(2)}(\hat{\mathbf{T}})=\{\hat{\mathbf{T}}_{i},1\leq i\leq N\}$
$g_{2}^{(1)}(\hat{\mathbf{H}})=\{\hat{\mathbf{H}}_{i},1\leq i\leq N\}$ $,g_{2}^{(2)}(\hat{\mathbf{T}})=\{\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\hat{\mathbf{T}}_{j},1\leq i\leq N\}$
(5) $g_{3}^{(1)}(\hat{\mathbf{H}})=\{\hat{\mathbf{T}}_{i},1\leq i\leq N\}$ $,g_{3}^{(2)}(\hat{\mathbf{T}})=\{\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\hat{\mathbf{T}}_{j},1\leq i\leq N\}$
where $\hat{\mathbf{H}}\in\mathbb{R}^{N\times d}$ denotes graph representations and $\hat{\mathbf{T}}\in\mathbb{R}^{N\times d}$ indicates text representations, and $N$ is the number of nodes.
A.4. Instruction Templates across Tasks/Data
We provide instruction templates for different tasks (graph matching, node classification, and link prediction) on various datasets in our experiment. The templates are shown in Table 8 and 9. While the same template is used for paper citation graphs, we design specific templates for each data in the downstream tasks, including both standard and COT (Chain-of-Thought) instructions. The <graph> token serves as an indicator, and ... denotes fillable blanks.
A.5. Supplementary Experimental Results
In this subsection, we analyze the link prediction performance using different instruction mixtures. We then compare the prediction results of our GraphGPT and ChatGPT under different instruction formats, including node content only, node content with text-based graph structure, and our custom-designed graph instruction.
A.5.1. Performance of Link Prediction
Four different instruction mixtures are compared: i) COT instruction data for node classification only (w/o Link) ii) Instruction data for link prediction only (only Link) iii) A mix of standard instruction data for node classification and instruction data for link prediction (Arxiv-std + PubMed-std + Link) iv) A mix of 50% standard instruction for node classification, 50% COT instruction data for node classification, and instruction for link prediction (Arxiv-mix + PubMed-mix + Link). When compared to representative baselines, as presented in Table 7, the combination of standard instruction data for node classification and instruction for link prediction significantly outperformed the baselines. Furthermore, as mentioned earlier in Table 2, this particular variant outperformed state-of-the-art approaches in node classification. This improvement can be attributed to the task of link prediction, which enhances our GraphGPT’s understanding of graph structural information and facilitates node classification tuning. Importantly, the inclusion of node classification instructions does not impede the model’s comprehension of link prediction.
A.5.2. Comparison Between ChatGPT and GraphGPT
Table 10 shows more comparison cases between ChatGPT and GraphGPT.

### v3

@@ 4.1.4. Implementation Details
For our model implementation, we primarily use the PyTorch and Transformers libraries. We utilize Vicuna-7B-v1.1 and Vicuna-7B-v1.5 as the base models. The batch size is set to 2 per GPU, and the learning rate is $2e^{-3}$. We apply a warmup ratio of $3e^{-2}$ and set the maximum input length of the Large Language Model (LLM) to 2048. The training process runs for 3 epochs. In the task-specific instruction tuning stage, we explore various combinations of instruction data to assess the model’s performance under different data mixtures. The hyperparameter settings remain constant, except for the number of training epochs, which is set to 2 in this stage. The alignment projector parameters fine-tuned in the self-supervised instruction tuning stage serve as the initial parameters for the projector in the second tuning stage. For evaluating most baselines, we use their publicly available code. For more implementation details, please refer to our released code.
@@ 4.2. Overall Performance Comparison (RQ1)
We conduct experiments on the node classification task, evaluating both supervised and zero-shot scenarios. The overall performance is summarized in Table 1. In the Supervised Task Setting, models are trained on a specific dataset and evaluated on the corresponding test set (e.g., training on Arxiv-Arxiv and testing on the Arxiv test set). In the Zero-Shot Task Setting, models are trained on a specific dataset and tested on other datasets without additional training (e.g., training on Arxiv-PubMed and testing on the PubMed dataset). To handle variations in the number of classes across datasets, we employ a transfer-trained classifier, such as a linear layer, when testing GNN-based models. In Table 1, ”-7B-” indicates the parameter scale, while ”-v1.1-” and ”-v1.5-” represent different versions of the base Vicuna model. ”-stage2” indicates the adoption of only the second stage tuning. ”-std” and ”-cot” denote the use of the standard and generated COT instruction datasets, respectively.
@@
LLM-based solutions like Baichuan-7B and Vicuna-7B maintain stable performance across different datasets but rely solely on text information for predictions. In contrast, our GraphGPT preserves graph structure, providing a comprehensive solution for graph learning tasks. Two key factors contribute to these improvements: (i) Our dual-stage graph instruction tuning aligns structural information encoded by the graph encoder with natural language tokens, enabling the LLM to understand the graph’s inherent characteristics. (ii) Our framework facilitates mutual enhancement between the graph encoder and LLM, filling the LLM’s gap in structural understanding and enabling it to reason about the graph’s structure.
@@
Table 3. Performance comparison of various instruction mixtures for link prediction on PubMed.
Dataset PubMed
Model AUC AP
MLP 0.5583 0.5833
GAT 0.5606 0.6373
GraphSAGE 0.5041 0.5813
RevGNN 0.4538 0.5083
Node2Vec 0.6535 0.6885
w/o Link 0.5010 0.5005
only Link 0.6704 0.6087
Arxiv-std + PubMed-std + Link 0.8246 0.8026
Arxiv-mix + PubMed-mix + Link 0.6451 0.5886
@@ 4.3. Generalization Ability Investigation (RQ2)
Generalization for Multitasking Graph Learner. Recent studies on instruction tuning suggest that mixing different instruction tuning data can further enhance the performance of Language and Logic Models (LLMs). In this study, we ensure a consistent number of instruction entries and mix different types of instruction data, including standard instruction (”-std”), COT instruction (”-cot”), a blend of standard (50%) and COT (50%) instruction (”-mix”), and link prediction instruction (”Link”). The results are presented in Tables 2 and Table 3. We can observe that effective data mixture solutions can significantly improve the performance of our GraphGPT under various settings. The addition of task-specific instruction for link prediction task notably enhances the performance of our model in node classification. Interestingly, after incorporating node classification, the performance of link prediction also exceeds that of the selected best-performing existing models. After mixing the instructions of different tasks, our model demonstrates the ability to effectively handle various graph learning tasks and transfer its knowledge to other unseen datasets.
@@ 4.6. Model Case Study (RQ5)
We conduct a detailed analysis of our model’s performance in downstream graph learning tasks compared to traditional LLMs using different types of instructions. We evaluate ChatGPT and our GraphGPT using Arxiv data, with prompts based on node content, node content with text-based graph structure, and our designed graph instruction. The results, shown in Table 6, clearly demonstrate that despite its massive parameter count (over 200B), ChatGPT struggles to make accurate predictions solely based on node text information or node content with text-based graph structure. This challenge is particularly evident when dealing with papers that have cross-disciplinary characteristics, as seen in the example of research domains in machine learning and hardware architecture. In contrast, our GraphGPT consistently provides accurate predictions and reasonable explanations. This is because our GraphGPT incorporates a subgraph structure with 103 nodes, allowing it to extract rich structural information from neighboring nodes’ citation relationships, leading to accurate predictions.
@@ 5. Related Work
Self-supervised Learning and Pre-training on Graphs. To enhance the robustness of graph models, self-supervised learning (SSL) has been introduced as a powerful technique (Liu et al., 2022; Jing et al., 2021; Hu et al., 2020b). It allows GNNs to learn meaningful graph representations without heavily relying on labeled data. The core idea behind self-supervised learning in graph models is to design pretext tasks that leverage the graph’s intrinsic properties to generate additional supervision signals (Xia et al., 2022). SSL-enhanced graph learning methods can be broadly classified into two main paradigms: contrastive SSL and generative SSL. In particular, i) Contrastive SSL focuses on learning representations by contrasting positive and negative samples. Notable methods in this domain include GraphCL (You et al., 2020) and GCA (Zhu et al., 2021). Recent advancements in contrastive SSL include automated contrastive augmentation (i.e., JOAO (You et al., 2021), AdaGCL (Jiang et al., 2023)), Hyperbolic-Euclidean dual space contrasting (e.g., DSGC (Yang et al., 2022a)), or community-aware contrastive learning (e.g., gCooL (Li et al., 2022)). ii) Generative SSL, on the other hand, focuses on generating realistic samples that resemble the original graph structures. Recent advancements in this line include GraphMAE (Hou et al., 2022, 2023) for feature masking, and S2GAE (Tan et al., 2023), AutoCF (Xia et al., 2023) for reconstructing masked edges as SSL tasks.